OpenStudy (anonymous):

if m=4, find a value for k such that y-intercept will be -2 y-2=m(x-k)

OpenStudy (anonymous):

first off, if \(m=4\) you have \[y-2=4(x-k)\] right?

OpenStudy (anonymous):

yeah. then i distributed 4 to x and k so i got y-2=4x-4k

OpenStudy (anonymous):

ok good now add \(2\) to both sides

OpenStudy (anonymous):

y=4x-4k+2 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which looks like \(y=mx+b\) where \(m=4\) and \(b=-4k+2\)

OpenStudy (anonymous):

so your job is to set \[-4k+2=-2\] and solve for \(k\) which takes two steps only

OpenStudy (anonymous):

okay because thats where i stopped and didnt know what to do. k=1

OpenStudy (anonymous):

thanks again